OpenStudy (anonymous):

simplify: sqrt(x^3(y^6(z^-5)))^3/(x^8(y^5)(z^4))

jhonyy9 (jhonyy9):

\[\sqrt{((x ^{9}y ^{18})/z ^{15})/(x ^{8}y ^{5}z ^{4})} =\]

jhonyy9 (jhonyy9):

\[\sqrt{(x ^{9}y ^{18}/x ^{8}y ^{5}z ^{19}} =\]

jhonyy9 (jhonyy9):

\[\sqrt{(xy ^{13})/z ^{19}} =\sqrt{(xy ^{12}y)/(z ^{18}z)} =(y ^{6}/z ^{9})\sqrt{(xy/z)}\]